# Quantifying the Importance of Data Alignment in Downstream Model Performance

### v2

@@ Abstract
Keywords: LLMs, large language model, Autoformalization, data centric machine learning, fine-tuning, automated reasoning
@@ 1 Introduction
Research within the domain of Large Language Models (LLMs) has historically placed an emphasis on the size of datasets used for pre-training, claiming it is one of the primary determinants of LLM performance Chowdhery et al. (2022); Nostalgebraist (2022); OpenAI (2023); Google (2023b). Empirical evidence demonstrates this trend, as models trained on large datasets exhibit superior performance. Notably, GPT-4, with its conjectured 1 petabyte dataset, markedly surpasses GPT-3—which is trained on a comparatively modest 45 terabytes—in terms of response quality and contextual accuracy OpenAI (2023). However, emerging research indicates that other dimensions, such as dataset diversity, play a crucial role in the efficacy of LLMs, with high-performing models often arising from datasets with high diversity coefficients Lee et al. (2023).
Current discourse predominantly highlights the scale of a dataset as a pivotal factor in its capacity to effectively pre-train or fine-tune a model, with emphasis frequently placed on quantitative metrics—specifically, the sheer size of the dataset. Lee et al. (2023) This investigation, however, seeks to shift this paradigm to consider qualitative assessments, notably the alignment of datasets with the specific evaluation tasks. Building upon methodologies established in previous studies for quantifying dataset alignment, our research aims to examine the role of data quality in the pre-training and fine-tuning process, verifying the hypothesis that increased data alignment could significantly improve LLM performance. This paradigm challenges the emphasis on dataset size, suggesting an alternative approach to dataset importance and optimization in the context of LLM training – i.e., select the most aligned data to your target task. We explore this via Autoformalization.
Autoformalization is defined as the transformation of concepts in natural language to formalized, structured language like mathematical proofs or code. The creation of a proficient Autoformalization tool would not only drastically reduce the substantial costs associated with manual formalization efforts but could also serve as a bridge linking the automated theorem verification and computational algebra with the extensive body of mathematical knowledge predominantly recorded in natural language. Moreover, the capacity for Autoformalization underscores a machine’s adeptness at navigating the subtleties of human language and the precision required by formal linguistic systems Wu et al. (2022).
@@ 2.2 Dataset Alignment Quantification
To quantify dataset alignment, we employ the Task2Vec Alignment Coefficient, which facilitates a rigorous comparative assessment of dataset similarity Lee et al. (2023).
@@ 3.2.1 Experimental Setup and Motivation
    AF Dataset (AF): A dataset consisting of informal statements and their formal counterparts in Isabelle designed for training LLMs to perform Autoformalization. We use its test set as a benchmark of LLM performance on statement Autoformalization. Thus, we also believe it will result in the lowest perplexity among the proof datasets when used to train an LLM for AF Miranda (2021).
@@
    The Stack Smol Python Docstrings dataset (Docstring): a dataset consisting of concise function headers written in informal language and their implementations in python; we use it to assess how well coding datasets can finetune for Autoformalization Bird (2023a).
@@
    The Stack Dedup Python Docstrings 1.0 percent unified dataset (Docstring 2): A dataset consisting of function headers written in informal language and their implementations in python; given its nature we anticipate it scoring among the lowest of perplexity scores against the Docstring benchmark Bird (2023b).
@@
    C4-EN-10K Dataset (C4): A ten-thousand-entry subset of a database composed of text pulled from Common Crawl (an internet archive) meant for pre-training for general English language modeling. Given it’s entries are all informal statements not related to mathematics, we predict a high perplexity score in performing AF Raffel et al. (2019).
@@
    wikitext-2-raw-v1 Dataset (Wikitext): A subset of the Wikitext dataset; Wikitext is a dataset composed of text taken from Wikipedia pages that met the score guidelines to qualify as either a ’good’ or ’featured’ article; given its nature and lack of relevance to AF, we expect a high perplexity score Merity et al. (2016).
@@
    minif2f-lean4 Dataset (LeanDojo4): A subset of the miniF2F dataset which is comprised of math exercise statements and their formal counterparts in lean; given that it is in a different formal language, we expect a mid-range perplexity score Zheng et al. (2021).
@@
    Proofnet Dataset (Proofnet): This dataset is comprised of statements taken from undergraduate math courses and their formal counterparts in lean; given their similarities, we expect LeanDojo4 and Proofnet to score similarly in perplexity Azerbayev et al. (2023).
@@
    HumanEvalPack: This dataset consists of a prompt describing a function and implementations of the function in Python, JavaScript, Java, Go, C++, and Rust as well as buggy solutions to serve as bad examples. We expect it to obtain a mid-range score against the Docstring benchmark Muennighoff et al. (2023).
@@ 3.3.2 Code Dataset Results
For our code dataset, we found again a strong negative correlation between alignment score and perplexity loss. Immediately we see that perplexity scores are much lower for the code datasets than the proof datasets. This is probably due to the fact that GPT-2 knows how to generate code quite well based solely on pre-training, as it has been pre-trained on a large, diverse web-based pre-training corpora Radford et al. (2019). As a result, fine-tuning further on code produces even greater results. Standard GPT-2 has a baseline perplexity score of  14.8, which is quite good and is indicated by the dotted gray line in Figure 2.
@@ 3.4 Impact of Data Alignment versus Dataset Size on LLM Performance
Broader Impact Statement
Validating our hypothesis, indicating a consistent improvement in Autoformalization results through data alignment, would have significant implications for the NLP field, particularly given the recent surge in attention towards LLMs. By demonstrating the critical role of data quality and alignment in optimizing LLM performance Lee et al. (2023), our work challenges the prevailing emphasis on dataset size as the principal metric for pre-training efficacy. Furthermore, our findings advocate for a more strategic allocation of computational resources, potentially leading to substantial time and cost savings in the development of LLMs across both academic and industry spheres.
Acknowledgments and Disclosure of Funding
We would like to acknowledge Michael Bernstein and the entire Stanford CS197 course staff for their valuable support and guidance throughout this project. In addition, we are grateful for being granted access to and use of the Stanford SNAP cluster GPU servers and the Stanford Trustworthy AI Research (STAIR) Lab’s compute resources for running our experiments. Brando Miranda, acknowledges support by NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, OpenAI and Google Inc. Brando Miranda acknowledges the Stanford Edge fellowship.
@@ F.1 Discussion of AF-Split Experiment Outcomes
The article Google (2023a) “PaLM 2 Technical Report” by Google discusses the development and performance of PaLM 2. The study showcases PaLM 2’s versatility but also emphasizes the role of architectural enhancements and diverse model objectives in achieving superior results. The inclusion of a diverse data mixture, even incorporating a small amount of translation pairs, results in performance comparable to dedicated translation services, a statement which supports our belief that data quality can be a critical factor in determining how well a dataset can train an LLM. This sentiment is also expressed in the article “Model Performance Scaling with Multiple Data Sources” by Tatsunori Hashimoto.Hashimoto (2021) It discusses the challenges of training ML models using data from various sources that vary in quality and cost. Hashimoto proposes a parametric model to approximate generalization error, which is more accurate for various models compared to existing linear approximations. The work represents a step toward better understanding model performance under varying data conditions and questions whether the approach can scale to more extreme scenarios or larger numbers of data sources in future research.

### v3

@@ Abstract
LLMs, large language model, Autoformalization, data centric machine learning, fine-tuning, automated reasoning
@@ 1 Introduction
Research within the domain of Large Language Models (LLMs) has historically placed an emphasis on the size of datasets used for pre-training, claiming it is one of the primary determinants of LLM performance (Chowdhery et al., 2022; Nostalgebraist, 2022; OpenAI, 2023; Google, 2023b). Empirical evidence demonstrates this trend, as models trained on large datasets exhibit superior performance. Notably, GPT-4, with its conjectured 1 petabyte dataset, markedly surpasses GPT-3—which is trained on a comparatively modest 45 terabytes—in terms of response quality and contextual accuracy (OpenAI, 2023). However, emerging research indicates that other dimensions, such as dataset diversity, play a crucial role in the efficacy of LLMs, with high-performing models often arising from datasets with high diversity coefficients (Lee et al., 2023).
Current discourse predominantly highlights the scale of a dataset as a pivotal factor in its capacity to effectively pre-train or fine-tune a model, with emphasis frequently placed on quantitative metrics—specifically, the sheer size of the dataset. (Lee et al., 2023) This investigation, however, seeks to shift this paradigm to consider qualitative assessments, notably the alignment of datasets with the specific evaluation tasks. Building upon methodologies established in previous studies for quantifying dataset alignment, our research aims to examine the role of data quality in the pre-training and fine-tuning process, verifying the hypothesis that increased data alignment could significantly improve LLM performance. This paradigm challenges the emphasis on dataset size, suggesting an alternative approach to dataset importance and optimization in the context of LLM training – i.e., select the most aligned data to your target task. We explore this via Autoformalization.
Autoformalization is defined as the transformation of concepts in natural language to formalized, structured language like mathematical proofs or code. The creation of a proficient Autoformalization tool would not only drastically reduce the substantial costs associated with manual formalization efforts but could also serve as a bridge linking the automated theorem verification and computational algebra with the extensive body of mathematical knowledge predominantly recorded in natural language. Moreover, the capacity for Autoformalization underscores a machine’s adeptness at navigating the subtleties of human language and the precision required by formal linguistic systems (Wu et al., 2022).
@@ 2.2 Dataset Alignment Quantification
To quantify dataset alignment, we employ the Task2Vec Alignment Coefficient, which facilitates a rigorous comparative assessment of dataset similarity (Lee et al., 2023).
@@ 3.2.1 Experimental Setup and Motivation
    AF Dataset (AF): A dataset consisting of informal statements and their formal counterparts in Isabelle designed for training LLMs to perform Autoformalization. We use its test set as a benchmark of LLM performance on statement Autoformalization. Thus, we also believe it will result in the lowest perplexity among the proof datasets when used to train an LLM for AF (Miranda, 2021).
@@
    The Stack Smol Python Docstrings dataset (Docstring): a dataset consisting of concise function headers written in informal language and their implementations in python; we use it to assess how well coding datasets can finetune for Autoformalization (Bird, 2023a).
@@
    The Stack Dedup Python Docstrings 1.0 percent unified dataset (Docstring 2): A dataset consisting of function headers written in informal language and their implementations in python; given its nature we anticipate it scoring among the lowest of perplexity scores against the Docstring benchmark (Bird, 2023b).
@@
    C4-EN-10K Dataset (C4): A ten-thousand-entry subset of a database composed of text pulled from Common Crawl (an internet archive) meant for pre-training for general English language modeling. Given it’s entries are all informal statements not related to mathematics, we predict a high perplexity score in performing AF (Raffel et al., 2019).
@@
    wikitext-2-raw-v1 Dataset (Wikitext): A subset of the Wikitext dataset; Wikitext is a dataset composed of text taken from Wikipedia pages that met the score guidelines to qualify as either a ’good’ or ’featured’ article; given its nature and lack of relevance to AF, we expect a high perplexity score (Merity et al., 2016).
@@
    minif2f-lean4 Dataset (LeanDojo4): A subset of the miniF2F dataset which is comprised of math exercise statements and their formal counterparts in lean; given that it is in a different formal language, we expect a mid-range perplexity score (Zheng et al., 2021).
@@
    Proofnet Dataset (Proofnet): This dataset is comprised of statements taken from undergraduate math courses and their formal counterparts in lean; given their similarities, we expect LeanDojo4 and Proofnet to score similarly in perplexity (Azerbayev et al., 2023).
@@
    HumanEvalPack: This dataset consists of a prompt describing a function and implementations of the function in Python, JavaScript, Java, Go, C++, and Rust as well as buggy solutions to serve as bad examples. We expect it to obtain a mid-range score against the Docstring benchmark (Muennighoff et al., 2023).
@@ 3.3.2 Code Dataset Results
For our code dataset, we found again a strong negative correlation between alignment score and perplexity loss. Immediately we see that perplexity scores are much lower for the code datasets than the proof datasets. This is probably due to the fact that GPT-2 knows how to generate code quite well based solely on pre-training, as it has been pre-trained on a large, diverse web-based pre-training corpora (Radford et al., 2019). As a result, fine-tuning further on code produces even greater results. Standard GPT-2 has a baseline perplexity score of  14.8, which is quite good and is indicated by the dotted gray line in Figure 2.
@@ 3.4 Impact of Data Alignment versus Dataset Size on LLM Performance
Impact Statement
Validating our hypothesis, indicating a consistent improvement in Autoformalization results through data alignment, would have significant implications for the NLP field, particularly given the recent surge in attention towards LLMs. By demonstrating the critical role of data quality and alignment in optimizing LLM performance (Lee et al., 2023), our work challenges the prevailing emphasis on dataset size as the principal metric for pre-training efficacy. Furthermore, our findings advocate for a more strategic allocation of computational resources, potentially leading to substantial time and cost savings in the development of LLMs across both academic and industry spheres.
@@ F.1 Discussion of AF-Split Experiment Outcomes
The article (Google, 2023a) “PaLM 2 Technical Report” by Google discusses the development and performance of PaLM 2. The study showcases PaLM 2’s versatility but also emphasizes the role of architectural enhancements and diverse model objectives in achieving superior results. The inclusion of a diverse data mixture, even incorporating a small amount of translation pairs, results in performance comparable to dedicated translation services, a statement which supports our belief that data quality can be a critical factor in determining how well a dataset can train an LLM. This sentiment is also expressed in the article “Model Performance Scaling with Multiple Data Sources” by Tatsunori Hashimoto.(Hashimoto, 2021) It discusses the challenges of training ML models using data from various sources that vary in quality and cost. Hashimoto proposes a parametric model to approximate generalization error, which is more accurate for various models compared to existing linear approximations. The work represents a step toward better understanding model performance under varying data conditions and questions whether the approach can scale to more extreme scenarios or larger numbers of data sources in future research.